OpenStudy (anonymous):

What is 0.123456712345671234567... (repeating) in fraction form.

OpenStudy (bahrom7893):

12345671234567(repeating)/10000000000000000...... lol

OpenStudy (anonymous):

lol

OpenStudy (bahrom7893):

Okay.. suppose x = 0.1234567... then 10,000,000x = 1234567.1234567...

OpenStudy (anonymous):

1234567/9999999

OpenStudy (bahrom7893):

10,000,000x-x=1234567 9,999,999x = 1234567 or x = 1234567/9,999,999

OpenStudy (anonymous):

right!

OpenStudy (youssef):

hi all

OpenStudy (2bornot2b):

Hello youssef

OpenStudy (youssef):

hello would you please help me in implicit differentiation

OpenStudy (2bornot2b):

Seems like you are new to OpenStudy... Do you want to post a question?

OpenStudy (2bornot2b):

I mean posting your question on implicit differentiation can help you a lot

OpenStudy (youssef):

\[\sqrt{y} lny+siny=1 find dy/dx\]

OpenStudy (2bornot2b):

youssef do you know that you are not being heard by the talented users of this site. Please post your question using the top left box labelled as "ask a question"

OpenStudy (2bornot2b):

Do you understand me?